# Instruction-Tuning LLMs for Event Extraction with Annotation Guidelines

### v1

@@ Abstract
In this work, we study the effect of annotation guidelines—textual descriptions of event types and arguments, when instruction-tuning large language models for event extraction. We conducted a series of experiments with both human-provided and machine-generated guidelines in both full- and low-data settings. Our results demonstrate the promise of annotation guidelines when there is a decent amount of training data and highlight its effectiveness in improving cross-schema generalization and low-frequency event-type performance. [^1] Our source code and datasets are available at https://github.com/salokr/LLaMA-Events.
@@ 1 Introduction
Finally, we confirmed a consistent effect of annotation guidelines on the smaller LLaMA-3.2-1B-Instruct model and conducted an in-depth analysis showing that the guidelines can help LLMs reduce common types of EE errors and are beneficial to event types with any frequency in the training set.
@@ 2.2 Instruction-Tuning LLMs with Annotation Guidelines
To develop a scalable and cost-effective approach for guideline generation, we employ a reverse engineering strategy, leveraging both annotated event examples and the strong generative capabilities of LLMs. As illustrated in Figure X, we construct a guideline generation prompt for each event type $e$ by providing a few annotated examples $\{(X_{i},Y_{i})\}$ demonstrating the existence or non-existence of event instance of type $e$, and then prompt an LLM (GPT-4o in our experiment) to generate annotation guidelines for $e$. In total, we experimented with five variants of machine-generated guidelines: (1) Guideline-P: We prompt the LLM with 10 positive annotated examples of type $e$ to generate the annotation guidelines. Inspired by Sainz et al. (2024), we sample 5 distinct guidelines for each event type, which can be used during the model training to ensure that the model is exposed to multiple rephrasings of the guidelines rather than memorizing and overfitting to a specific one. (2) Guideline-PN (Positive + Negative Examples): In addition to 10 positive event annotations, we also provide 15 negative annotations where the input $X$ does not imply event instances of type $e$. Similarly, we prompt the LLM to generate 5 distinct guidelines for each event type. (3) Guideline-PS (Positive + Sibling Events): Similar to Guideline-PN, we prompt the LLM with both positive and negative event annotations. However, the negative annotations are selected from the sibling event types of the target type $e$ (e.g., Arrest vs. Jail), as defined by the event ontology. We hypothesize that the critical challenge for EE lies in distinguishing between sibling event types; hence, an instructed LLM can benefit from following annotation guidelines that particularly emphasize the difference between sibling event types. As in the earlier variants, we generate 5 guideline samples per event type. (4) Guideline-PN-Int and (5) Guideline-PS-Int: Finally, we create two more variants that Integrate the 5 diverse guideline samples from Guideline-PN and Guideline-PS into a comprehensive one, respectively. Examples of the 5 guideline variants are shown in Table 1. The prompt templates used for generating guidelines and example generations are provided in Appendix C and D, respectively.
@@ 3.1 Experimental Setup
We include all details about datasets, evaluation, and model training in Appendix A-B. We will open-source our scripts for automatically converting datasets in TextEE Huang et al. (2024) into Python code format (we dub the processed version as PyCode-TextEE) and for evaluating extracted events automatically via code execution at our GitHub repository https://github.com/salokr/LLaMA-Events.
@@ 3.2 RQ1: Do the annotation guidelines allow an LLM to more precisely extract occurring events?
To assess the impact of incorporating annotation guidelines in the EE instructions, we compare instruction-tuning an LLM with and without guidelines. We hypothesize that including the annotation guidelines can help the LLM more easily distinguish between similar event types. To understand its impact, we also compare this approach with a “negative sampling (NS)” approach. Specifically, we instruction-tune the LLM on an augmented training set, where each training example is supplemented with 15 randomly selected negative samples, i.e., triplets of $(e_{neg},X,\phi)$ with non-existing event type $e_{neg}$ yielding empty extraction output. We note that annotation guidelines and negative sampling are two complementary approaches for an LLM to learn to distinguish between event types. In our experiments, we thus evaluated the effect of annotation guidelines in two independent settings: (1) training on the original training set (w/o NS) and (2) training on the negative sample-augmented training set (w/ NS).
@@ 3.6 Further Analysis
As shown in Figure 3, frequent event types show consistent improvements with annotation guidelines, as indicated by the green bars, suggesting that even well-represented events benefit from enhanced annotations. Only a few declines (red bars) occurred with mid- to low-frequency event types, whereas most event types still benefit from the guidelines. In fact, these event types, especially the very rare ones (top of the figure), generally present larger gains (i.e., longer green bars) than the more frequent types, which demonstrates a unique advantage of annotation guidelines in low-resource settings.
@@ 4 Related Work
With the growing capabilities of LLMs, recent efforts have explored their potential in IE Xu et al. (2024) and studied EE as an auxiliary task. Existing LLM-based IE methods generally fall into two categories: In-Context Learning (ICL) and Supervised Fine-Tuning (SFT). ICL-based approaches Li et al. (2023b); Guo et al. (2023); Ashok and Lipton (2023); Wang et al. (2023b) rely on providing a few-shot context within prompts, enabling LLMs to infer structured information without explicit parameter updates. While being data-efficient, they were found to misinterpret the task specifications Gao et al. (2024) and suffer from brittle sensitivity to prompt phrasing and example ordering (Gao et al., 2023). In addition, they also incur prohibitive costs due to the lengthy reasoning chains especially for complex tasks. In contrast, SFT-based methods Lu et al. (2023); Wang et al. (2023a); Gui et al. (2024); Zhou et al. (2024) fine-tune LLMs on annotated datasets, which can significantly improve their EE performance. Our work deepens this line of research and particularly explores the inclusion of annotation guidelines in instructions. While there have been existing works on similar topics, they did not focus on EE Sainz et al. (2024) or instruction tuning Pang et al. (2023).
@@ 6 Limitations
While our study demonstrates the benefits of structured annotation guidelines for event extraction, several limitations remain. First, our evaluation is limited to two datasets (ACE and RichERE), both within the news domain, which may not fully capture how guidelines generalize to other domains such as biomedical or legal text. Future work should assess whether schema differences in other domains exhibit similar trends. Second, while we analyze guideline length and diversity, we do not explicitly optimize guideline generation, leaving open the question of how to best balance conciseness and informativeness. Exploring adaptive methods that retrieve or refine guidelines dynamically during training and inference could further improve efficiency. Lastly, our study primarily focuses on instruction-tuning an LLM with predefined event schemas; however, real-world applications often require handling previously unseen event types. Investigating how structured guidelines can aid zero-shot or few-shot event extraction remains an important avenue for future research.
@@ Appendix A Preprocessing and Data Sampling
Appendix B Evaluation Methodology and Metrics
@@
Appendix C Prompt Design and Model Training
@@ 1. Integrated the 5 definitions under "Event Definition" into a single definition:
Appendix D Dataset Examples Across Multiple Guideline Settings
@@
Listing 3: Prompt example for generating consolidated guidelines: Guideline-PN-Int, and Guideline-PS-Int.

### v2

@@ Abstract
In this work, we study the effect of annotation guidelines—textual descriptions of event types and arguments, when instruction-tuning large language models for event extraction. We conducted a series of experiments with both human-provided and machine-generated guidelines in both full- and low-data settings. Our results demonstrate the promise of annotation guidelines when there is a decent amount of training data and highlight its effectiveness in improving cross-schema generalization and low-frequency event-type performance. [^1] Our source code and datasets are available at https://github.com/Ziyu-Yao-NLP-Lab/PyCode-TextEE.
@@ 1 Introduction
Finally, we validated the consistency of our findings across both model scale and diversity. Using the smaller LLaMA-3.2-1B-Instruct model, we observed that annotation guidelines retain their benefits, reducing common EE errors and supporting both frequent and rare event types. Extending beyond the LLaMA family, experiments with Qwen2.5-Coder-1.5B-Instruct Hui et al. (2024), a code-oriented model with a distinct pertaining objective, revealed similarly robust gains. We also evaluated on Speed++ Parekh et al. (2024), a domain-shifted dataset of informal epidemic-related tweets, where guidelines continued to significantly improve performance. Together, these results confirm the broad utility of machine-generated guidelines across model architectures, data regimes, and textual domains.
@@ 2.2 Instruction-Tuning LLMs with Annotation Guidelines
To develop a scalable and cost-effective approach for guideline generation, we employ a reverse engineering strategy, leveraging both annotated event examples and the strong generative capabilities of LLMs. As illustrated in Figure 1, we construct a guideline generation prompt for each event type $e$ by providing a few annotated examples $\{(X_{i},Y_{i})\}$ demonstrating the existence or non-existence of event instance of type $e$, and then prompt an LLM (GPT-4o in our experiment) to generate annotation guidelines for $e$. In total, we experimented with five variants of machine-generated guidelines: (1) Guideline-P: We prompt the LLM with 10 positive annotated examples of type $e$ to generate the annotation guidelines. Inspired by Sainz et al. (2024), we sample 5 distinct guidelines for each event type, which can be used during the model training to ensure that the model is exposed to multiple rephrasings of the guidelines rather than memorizing and overfitting to a specific one. (2) Guideline-PN (Positive + Negative Examples): In addition to 10 positive event annotations, we also provide 15 negative annotations where the input $X$ does not imply event instances of type $e$. Similarly, we prompt the LLM to generate 5 distinct guidelines for each event type. (3) Guideline-PS (Positive + Sibling Events): Similar to Guideline-PN, we prompt the LLM with both positive and negative event annotations. However, the negative annotations are selected from the sibling event types of the target type $e$ (e.g., Arrest vs. Jail), as defined by the event ontology. We hypothesize that the critical challenge for EE lies in distinguishing between sibling event types; hence, an instructed LLM can benefit from following annotation guidelines that particularly emphasize the difference between sibling event types. As in the earlier variants, we generate 5 guideline samples per event type. (4) Guideline-PN-Int and (5) Guideline-PS-Int: Finally, we create two more variants that Integrate the 5 diverse guideline samples from Guideline-PN and Guideline-PS into a comprehensive one, respectively. Examples of the 5 guideline variants are shown in Table 1. The prompt templates used for generating guidelines and example generations are provided in Appendix C and E, respectively.
@@ 3.1 Experimental Setup
We include all details about datasets, evaluation, and model training in Appendix A-C. We will open-source our scripts for automatically converting datasets in TextEE Huang et al. (2024) into Python code format (we dub the processed version as PyCode-TextEE) and for evaluating extracted events automatically via code execution at our GitHub repository https://github.com/Ziyu-Yao-NLP-Lab/PyCode-TextEE.
@@ 3.2 RQ1: Do the annotation guidelines allow an LLM to more precisely extract occurring events?
To assess the impact of incorporating annotation guidelines in the EE instructions, we compare instruction-tuning an LLM with and without guidelines. We hypothesize that including the annotation guidelines can help the LLM more easily distinguish between similar event types. To understand its impact, we also compare this approach with a “negative sampling (NS)” approach. Specifically, we instruction-tune the LLM on an augmented training set, where each training example is supplemented with 15 randomly selected negative samples, i.e., triplets of $(e_{neg},X,\phi)$ with non-existing event type $e_{neg}$ yielding empty extraction output. We note that annotation guidelines and negative sampling are two complementary approaches for an LLM to learn to distinguish between event types. In our experiments, we thus evaluated the effect of annotation guidelines in two independent settings: (1) training on the original training set (w/o NS) and (2) training on the negative sample-augmented training set (w/ NS). Additional details on example selections are provided in Appendix B.
@@ 3.6 Further Analysis
Setup TI TC AI AC
NoGuide w/o NS 33.91 33.91 25.24 23.82
PN w/o NS 32.38 32.38 26.19 25.32
NoGuide w/ NS 46.33 46.33 36.29 33.22
PN w/ NS 68.92 65.54 48.25 45.57
Table 7: RichERE results with Qwen2.5-Coder-1.5B. Machine-generated guidelines and negative samples consistently improve performance.
Generalization to a CodeLLM
To further assess the robustness and generalizability of our findings beyond the LLaMA family, we conducted additional experiments using Qwen2.5-Coder-1.5B-Instruct, a code-oriented LLM with a distinct architectural design and pretraining objective. We focused specifically on replicating the RichERE experiments (see Table 6), comparing the PN Guideline variant against the baseline NoGuideline, both with and without negative samples. As shown in Table 7, Qwen2.5-Coder demonstrates substantial improvements. Notably, we observe a 2% absolute increase in AC performance when guidelines are applied without negative samples (23.82 Vs. 25.32) and a significant 12.35% gain in TI performance when negative samples are included (46.33 Vs. 68.92). These consistent improvements further reinforce the efficacy of incorporating machine-generated annotation guidelines, highlighting their generalizability across diverse LLM architectures and instruction-tuning paradigms.
Setup TI TC AI AC
NoGuide w/o NS 31.99 30.56 22.27 21.57
PN w/o NS 43.03 43.03 29.19 28.15
NoGuide w/ NS 47.95 47.08 34.01 32.50
PN w/ NS 62.56 62.56 42.00 40.18
Table 8: Speed++ with LLaMA-3.2-1B. Machine-generated guidelines and negative samples consistently improve performance.
Generalization Beyond the News Domain
To study the utility of guidelines for domain generalization, we extended our evaluation to Speed++ Parekh et al. (2024), a dataset from the epidemic domain containing informal social media posts (tweets), contrasting significantly with the formal newswire text in ACE05 and RichERE. We utilized the same code-format conversion pipeline employed in previous experiments to maintain methodological consistency. For these experiments, we selected the LLaMA-3.2-1B model, aligning with our earlier analysis settings. Although our primary experiments leveraged larger models (e.g., LLaMA-3.1-8B), we opted for the smaller variant due to resource constraints. Nevertheless, this setup provides sufficient validation of our guideline approach across diverse textual domains. Table 8 summarizes our findings. Consistent with previous outcomes, we observed substantial improvements when incorporating machine-generated guidelines. Specifically, with negative sampling (w/ NS), TI increased by 14.61% (47.95 Vs. 62.56), and AC improved by 7.68% (32.50 Vs. 40.18). These results further underscore the robustness and generalizability of our proposed method, effectively extending its applicability beyond formal news articles to informal social media contexts.
@@
Figure 3 visualizes the change in AC scores across individual event types for both ACE05 and RichERE. It compares the performance of prompts without guidelines (dashed lines) against those enriched with machine-generated guidelines (solid lines). On average, machine-generated guidelines improve AC scores by +5.47 points on ACE05 (29.73% Vs. 35.2%) and +1.86 points on RichERE (25.32% Vs. 27.18%).
Event types on the y-axis are sorted by training frequency, and the green/red bars show per-event gains or drops. Among the 15 least frequent event types (indices $\geq$15), we observe gains in 8 ACE05 types and 5 RichERE types. For the remaining 5 event types in each dataset show no improvement (+0.0%) is observed, and a small decline is observed for 2 types in ACE05 and 5 in RichERE.
Upon inspection, most of the event types with limited or negative gains have extremely sparse training data, often less than 10 examples (e.g., Marry event), and in some cases as few as 2-3 instances (e.g., Acquit event). In such settings, the model may not receive enough signals to generalize. While our machine-generated guidelines cannot fully compensate for this sparsity, they still improve performance on several rare events, highlighting their potential to support low-resource settings.
@@ 4 Related Work
With the growing capabilities of LLMs, recent efforts have explored their potential in IE Xu et al. (2024) and studied EE as an auxiliary task. Existing LLM-based IE methods generally fall into two categories: In-Context Learning (ICL) and Supervised Fine-Tuning (SFT). ICL-based approaches Li et al. (2023b); Guo et al. (2023); Ashok and Lipton (2023); Wang et al. (2023b) rely on providing a few-shot context within prompts, enabling LLMs to infer structured information without explicit parameter updates. While being data-efficient, they were found to misinterpret the task specifications Gao et al. (2024) and suffer from brittle sensitivity to prompt phrasing and example ordering (Gao et al., 2023). In addition, they also incur prohibitive costs due to the lengthy reasoning chains especially for complex tasks. In contrast, SFT-based methods Lu et al. (2023); Wang et al. (2023a); Gui et al. (2024); Zhou et al. (2024); Wei et al. (2024) fine-tune LLMs on annotated datasets, which can significantly improve their EE performance. Our work deepens this line of research and particularly explores the inclusion of annotation guidelines in instructions. While there have been existing works on similar topics, they did not focus on EE Sainz et al. (2024) or instruction tuning Pang et al. (2023).
@@ 6 Limitations
While our study demonstrates the benefits of structured annotation guidelines for event extraction, several limitations remain. First, while our evaluation includes ACE and RichERE, both within the news domain, and an additional experiment on Speed++ from the epidemic domain, our study still lacks coverage of other domains such as biomedical or legal texts. This limited domain diversity may affect the generalizability of our findings. Future work should assess whether schema differences in other domains exhibit similar trends. Second, while we analyze guideline length and diversity, we do not explicitly optimize guideline generation, leaving open the question of how to best balance conciseness and informativeness. Exploring adaptive methods that retrieve or refine guidelines dynamically during training and inference could further improve efficiency. Lastly, our study primarily focuses on instruction-tuning an LLM with predefined event schemas; however, real-world applications often require handling previously unseen event types. Investigating how structured guidelines can aid zero-shot or few-shot event extraction remains an important avenue for future research. Additionally, future work could explore how machine-generated guidelines can assist human annotators during dataset creation, potentially improving annotation consistency and inter-annotator agreement in complex event extraction tasks.
@@ Appendix A Preprocessing and Data Sampling
Appendix B Selection of Negative Examples for Guideline Generation
For each event type, we selected 15 negative examples using a fixed random seed to ensure reproducibility. These negative examples were sampled from event instances belonging to types other than the target event type, as required by the definition of Guideline-PN. Each of these examples was used to inform the LLM that a given input does not express the event type in question. In our preliminary explorations, we did not observe any obvious discrepancy when we repeated the guideline generation with different random selections of 15 negative examples.
Appendix C Evaluation Methodology and Metrics
@@
Appendix D Prompt Design and Model Training
@@ 1. Integrated the 5 definitions under "Event Definition" into a single definition:
Appendix E Dataset Examples Across Multiple Guideline Settings
@@
Listing 3: Dataset example from ACE-05 with no guidelines.